OpenStudy (aravindg):

what us completing the square method..i knew it but frgt can anyone help me ?

OpenStudy (lgbasallote):

do you want a demonstration?

OpenStudy (aravindg):

general involving a ,b c and if possible a demonstration also

OpenStudy (aravindg):

it seems this method is pretty useful for solving quadratics thats why i am revising it

OpenStudy (lgbasallote):

firstly.... you agree that \[ax^2 + bx + c = 0\] that's the quadratic equation, yes?

OpenStudy (aravindg):

yep

OpenStudy (lgbasallote):

now..divide ALL terms by a...what do you get?

OpenStudy (lgbasallote):

where's the x?

OpenStudy (aravindg):

x^2+(b/a)x+(c/a)=0

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now subtract c/a from both sides

OpenStudy (aravindg):

\[x^2+(b/a)x=-c/a\]

OpenStudy (lgbasallote):

good. now divide b/a by 2..what do you get?

OpenStudy (aravindg):

only b/a?

OpenStudy (lgbasallote):

yes... but don't touch the equation yet

OpenStudy (lgbasallote):

just divide b/a by 2

OpenStudy (aravindg):

\[(x+b/2a)^2=(4c-b^2)/4a\]

OpenStudy (aravindg):

4c+b^2 srry

OpenStudy (lgbasallote):

why is it - b^2

OpenStudy (aravindg):

4c+b^2 srry

OpenStudy (lgbasallote):

\[x^2 + (\frac ba)x + \frac{b^2}{4a^2} = -\frac ca + \frac{b^2}{4a^2}\] \[\implies (x + \frac ba)^2 = \frac{b^2 - 4ac}{4a^2}\] you rushed too fast

OpenStudy (aravindg):

i see

OpenStudy (aravindg):

thx

OpenStudy (lgbasallote):

now take the square root of both sides

OpenStudy (lgbasallote):

do you need further help?

OpenStudy (aravindg):

nop gt it

OpenStudy (lgbasallote):

okay then

OpenStudy (aravindg):

@lgbasallote can you tell me how we can say that this method will be easier when we have an arbitrary quadratic?i mean is there any relation btw a,b,c so that i can recognise "AHA I CAN USE COMPLETING THE SQUARE HERE INSTEAD OF OTHER METHODS!"

OpenStudy (lgbasallote):

you can use completing the square method anywhere